The difference of two numbers is three. The larger number is 33 minus twice the smaller number. What are the numbers?
Answer by stanbon(75887)

The difference of two numbers is three. The larger number is 33 minus twice the smaller number. What are the numbers?
----------------------------------
Let the numbers be x and x+3.
EQUATION:
x+3 = 33 - 2x
3x = 30
x = 10
x+3 = 13
=====================
Cheers,
Stan H.